OpenStudy (anonymous):

What is the value of g[f(3)] for the functions f(x) = 2x + 3 and g(x) = x + 1? (Type only the value of your final answer, no variables.)

OpenStudy (anonymous):

\[f(3)=2\times 3+3=6+3=9\] \[g(9)=9+1=10\] so \[g(f(3))=10\]

OpenStudy (anonymous):

thanx